# Figure 1. Confirmation–Diagnosis–Correction–Redo (CDCR) Pattern. Based on the expected time of user interaction under the CDCR pattern, our model decides whether the agent should proceed autonomously or prompt the user to step in and supervise.:

### v1

@@ 8.2. Designing Confirmation as Interaction
To better balance user control with agent automation, confirmation should be treated not only as a scheduling problem but also as an interaction design challenge. From our study, we identify three insights that illustrate new ways of structuring this interaction.
8.2.1. Proactive engagement
In our two user studies, current agent UIs left users in a passive “wait state” during execution: with no meaningful hooks to monitor or steer the run, participants turned to other tasks and disengaged from supervision. This disengagement is risky and can lead to not only time and resources waste (discussed in our paper), but also harmful or unsafe actions (e.g., privacy breaches, opening phishing links, manipulative suggestions) (Alberts et al., 2024; Bansal et al., 2024; Amodei et al., 2016). Current systems like GPT Agents (OpenAI, 2025) exemplify a reactive model, asking for clarification only when they cannot proceed, which is insufficient to prevent harmful actions. As agents become increasingly autonomous, effective mixed-initiative interaction requires what Tennenhouse (2000) envisioned: moving humans from being “in the loop” to supervisory and policy-making roles. This shift—from human-centered to human-supervised design—requires agents to proactively solicit user supervision. Our model adds one dimension of proactivity by placing more checkpoints, but proactive supervision should extend beyond correctness to encompass broader forms of risk.
8.2.2. Goal Alignment
@@
Our findings on error location highlight the importance of prevention. Errors introduced early often stem from poorly communicated goals rather than faulty execution (Bansal et al., 2024). While most HCI work emphasizes co-execution or error repair (Feng et al., 2024; Arawjo et al., 2024; Epperson et al., 2025), we identify a design opportunity in early-stage goal alignment. Before execution, agents could present a structured plan for user review, drawing on concepts from formal languages that use rigorous specifications to evaluate whether task goals are satisfied (Aeronautiques et al., 1998; Shivashankar et al., 2013; Nau et al., 2003). As noted in Section 7.2.3, our model assigns each agent action its own correctness probability $p_{a_{i}}$, which could be adapted to support such early-stage goal alignment interfaces, enabling more frequent and dynamic checkpoints.
8.2.3. Right amount of information
As tasks grow longer, participants in our study found confirmation increasingly difficult: extended linear logs made it hard to locate issues, leaving users disoriented. This highlights a core design question—what level of detail should agents expose for effective supervision? Current UIs that present step-by-step activity lists with screenshots are often too low-level, overwhelming users with detail without offering structure. Future interfaces should balance informativeness and abstraction, allowing users to monitor actions without being lost in execution traces. One promising direction is to summarize and restructure agent actions hierarchically, so that users can reason about plans at multiple levels of abstraction rather than scanning long linear logs. If such hierarchical structuring is adopted, our CDCR pattern remains intact, but diagnosis time may decrease in a logarithmic fashion—since users could locate errors by checking higher-level chunks first (e.g., binary or ternary splits) rather than scanning all low-level steps.

### v2

@@ 8.2. Designing Confirmation for Transparency
To better balance user control with agent automation, confirmation should be treated not only as a scheduling problem but also as an interaction design challenge. We focus on three aspects: goal alignment, information exposure, and integration with downstream agents for transparency.
8.2.1. Plan alignment and control hooks
@@
Our findings on error location highlight the importance of prevention. Errors introduced early often stem from poorly communicated goals rather than faulty execution (Bansal et al., 2024). While most HCI work emphasizes co-execution or error repair (Feng et al., 2024; Arawjo et al., 2024; Epperson et al., 2025), we identify a design opportunity in early-stage goal alignment. Before execution, agents could present a structured plan for user review, drawing on concepts from formal languages that use rigorous specifications to evaluate whether task goals are satisfied (Aeronautiques et al., 1998; Shivashankar et al., 2013; Nau et al., 2003). Recent agentic interfaces already move in this direction by exposing intermediate plans and approval guards to users, allowing them to correct goals, such as Magentic-UI and Ongoal (Mozannar et al., 2025; Coscia et al., 2025). As noted in Section 7.2.3, our model assigns each agent action its own correctness probability $p_{a_{i}}$, which can be used to decide which parts of a proposed plan warrant mandatory review or additional clarification. In other words, these downstream agents define the interaction hooks for goal alignment, while our scheduler provides a way to time and target those hooks on the most consequential steps.
8.2.2. Targeted transparency and information granularity
As tasks grow longer, participants in our study found confirmation increasingly difficult: extended linear logs made it hard to locate issues, leaving users disoriented. This highlights a core design question—what level of detail should agents expose for effective supervision? Current UIs that present step-by-step activity lists with screenshots are often too low-level, overwhelming users with detail without offering structure. Recent systems surface the same tension in different domains: some agents selectively pause at ambiguous decision points (Peng et al., 2025; Feng et al., 2024; Huq et al., 2025), some agents visualize execution traces in hierarchical way (Zhou and MacLellan, 2024), and image agents maintain explicit intent states or reuse prior user preference models to reduce the number of refinement rounds (Hahn et al., 2025; Li et al., 2025b). All of these systems must implicitly decide how much internal state and history to show, and when asking for additional input is worth the user’s effort. Agents should offer targeted transparency, surfacing additional detail, explanations, or plan summaries only at checkpoints where the expected benefit of more information outweighs the cost of user attention. Hierarchical representations of agent behavior are one concrete instantiation of this idea: by grouping low-level actions into higher-level chunks, interfaces can let users quickly scan the structure of a run and then drill down only where something looks suspicious. If such hierarchical structuring is adopted, our CDCR pattern remains intact, but diagnosis time may decrease in a logarithmic fashion—since users could locate errors by checking higher-level chunks first (e.g., binary or ternary splits) rather than scanning all low-level steps. Our scheduler provides a normative basis for deciding where these targeted transparency checkpoints should sit. The same logic can be used to decide when to reveal more of the agent’s internal state as models become more accurate and routine errors become rare, shifting the role of confirmation from catching frequent mistakes to timing when additional transparency is most valuable.
@@ 9. Conclusion
Appendix A Formative Study Task Prompt
A.1. Task 1
Fill the cart on Kroger.com for a grocery delivery at 30332. I need all the ingredients to make a pizza at home. Maximize the quality of the ingredients while staying under the budget of 20 dollars.
Observed Errors: The agent mentioned adding certain items to the cart but those items were not present at the end of the run. The final summary from the model did not match the contents of the cart.
A.2. Task 2
Please help me edit this image. I want to create an advertisement for Coca Cola. Remove all the objects on the table except for the sprite. Change the sprite to coke. Remove the man in the background and change the background to an outdoor scene.
Observed Errors: The agent often applied wrong filters, selected objects improperly. The final result was very far from the desired result.
A.3. Task 3
Help me book a hotel in NYC for July 4 through July 12. It should be in Manhattan, clean, and within walking distance to several cafes. Look for good deals where the hotel is either less expensive than similar quality or on discount from its usual price.
Observed Errors: The agent often selected the wrong dates for flight/hotel bookings.
A.4. Task 4
Please turn these invoices into a spreadsheet which record each expense’s time and cost. And then send this spreadsheet to my email.
Observed Errors: There was a mismatch between the values provided in the invoice and the ones entered into the spreadsheet.
A.5. Task 5
Please generate a step-by-step plan for cooking onion soup In the Overcooked game, there are two primitive actions: 1. go to: ”goes to and faces the target object, where object is something like pot, onion, etc.” 2. interact: ”interact with the object, e.g., this should be called if you are trying to interact with the pot, onion dispenser, plate dispenser, tomato dispenser, tomato, onion, etc.” 3. wait20: waits for 20 time steps The objects in the environment are pot, onion dispenser, plate dispenser, tomato dispenser, tomato, onion. Rules: 1. you should have onion soup before you cook an onion 2. every time you cook something, you have to wait 20 minutes
Observed Errors: The agent selected the wrong location or did not execute the action it planned.
Appendix B Model Verification Study
B.1. Demographic Data
Demographic Information Participant Counts
AI Tool Usage Frequency Never (1%), Occasionally (13%), Weekly (22%), Daily (65%)
Age 18–25 (69%), 26–35 (30%), ¿55 (1%)
Gender Male (64%), Female (36%)
AI Tool Usage Proficiency Unfamiliar (16%), Beginner (18%), Intermediate (39%), Advanced (28%)
Table 1. Participant Information
(a) The Login Screen
(b) The Tutorial Screen.
Figure 9. User Interface Introduction
B.2. User Study Design
The user studies were conducted using a custom web application that presented participants with a simulated interface to an AI agent. The application was developed using standard HTML and CSS, with all experiment data logged to a Firebase Firestore database for subsequent analysis. The simulation comprised multiple screens that the user navigated sequentially to progress through the task.
Tutorial and Login : At the onset of the study, each user was presented with a short video tutorial within a modal popup, explaining the system’s UI components. The tutorial demonstrated a person using the identical interface to monitor a simulated agent during a task. The introduction also included annotated images highlighting key UI elements, such as the agent state display and control buttons. After closing the tutorial, users were asked to enter their participant ID, which was assigned during recruitment.
Task Execution: Upon logging in, users were presented with a task summary and a description of the expected outcome. Data logging was initiated when the user clicked the “Start Task” button, transitioning them to the “Working” page. This page displayed a loading spinner and the text “Agent is working” to simulate the processing latency of a real AI agent between confirmation points.
(a) Task Start Screen
(b) Working Screen
Figure 10. Main task interface
Experiment Design : Each participant ID was assigned a fixed sequence of experiments selected from a pool of image and tag pairs outlining the task flow. Each experiment was stored as a JSON object containing the step name, image reference, and a boolean value indicating correctness (used to determine which step contained an error). This sequence was structured to support the study’s within-subjects design.
B.3. User Study Task Description
Task 1: Shopping Cart Task : In this task, the user was presented with a simulated AI agent tasked with purchasing a predetermined list of items from a home delivery platform. The user monitored the agent’s progress and was asked to identify any errors the agent committed whenever a checkpoint was reached.
  • Subtasks:
    1. (1)
      Type the name of the item in the search bar and initiate the search.
    2. (2)
      Scroll through the results to identify the cheapest offering.
    3. (3)
      Add the selected offering to the cart.
Task 2: Image Editing Task : In this task, the user monitored an AI agent responsible for editing an image to achieve a specific target output. The input images contained several objects that required removal or replacement to generate the final result.
  • Subtasks:
    1. (1)
      Locate the object of interest within the frame.
    2. (2)
      Outline the object with high precision.
    3. (3)
      Remove or replace the object.
    4. (4)
      Repeat the steps above until the task is complete.
Task 3: Overcooked Task : This task involved monitoring an AI agent controlling a character within a simulated version of the game Overcooked. The agent was tasked with adding a specific list of ingredients to a pot and periodically adjusting the temperature to correctly cook the dish.
  • Subtasks:
    1. (1)
      Navigate to the item dispenser.
    2. (2)
      Retrieve the correct quantity of the specified item.
    3. (3)
      Add the item to the pot and adjust the temperature.
    4. (4)
      Repeat the previous steps for all listed ingredients.
Appendix C Simulation
C.0.1. Setup and Parameterization
We instantiated the image-editing task using the same base parameters as in our main simulation: a 16-step task with $p=0.91$, $t_{\text{confirm}}=10$, $t_{\text{diagnose}}=3$, and $t_{\text{redo}}=10$. To demonstrate that our model naturally supports heterogeneous, per-step specifications, we introduced controlled step-wise variations on top of this base regime.
C.1. Heterogeneous Step-Wise Parameters
C.1.1. Step-Wise Probabilities and an Error Cluster
We assigned non-uniform correctness probabilities across the 16 steps: the first five steps are moderately reliable ($p_{a_{1}}=\cdots=p_{a_{5}}=0.9$), the next four steps form a brittle error-prone cluster ($p_{a_{6}}=\cdots=p_{a_{9}}=0.75$), and the remaining seven steps are highly reliable ($p_{a_{10}}=\cdots=p_{a_{16}}=0.95$). As expected, the optimal checkpoint schedule increases its confirmation density within the brittle segment, indicating that the model adapts intervention frequency to localized spikes in error likelihood, as shown in Figure  11.
Figure 11. Step-wise correctness probabilities and the corresponding optimal checkpoint schedule
The first five steps have moderate reliability (\$p=0.9\$), steps 6–9 form an error-prone cluster (\$p=0.75\$), and steps 10–16 are highly reliable (\$p=0.95\$). The model allocates denser checkpoints within the brittle region (steps 4, 6, 7, 9, 10), and sparsifies them in the more reliable segments, adaptively concentrating user supervision where errors are most likely to arise.
C.1.2. Step-Wise Timing Parameters
To probe how heterogeneous timing affects the optimal schedule, we next made the confirmation time depend on the position of the step in the task. Specifically, for a 16-step task we set
$t_{\text{confirm}}^{k}=5+0.5\,(k-1),\quad k=1,\dots,16,$
so that confirming later steps becomes progressively more expensive. As a result, the optimal policy spreads checkpoints farther apart toward the end of the trajectory, trading off higher accumulated error risk against the rising marginal cost of confirmation, as shown in Figure  12.
Figure 12. Interval-dependent confirmation time and the resulting optimal checkpoint schedule
[…]
We additionally simulated step-wise variants of diagnosis time $t_{\text{diagnose}}^{k}$ and execution/redo time $t_{\text{redo}}^{k}$. Across these heterogeneous timing regimes, we observed consistent qualitative trends: increasing confirmation time leads to wider spacing between checkpoints, while decreasing diagnosis or execution time makes it cheaper to discover and repair errors after they occur, and thus also increases the optimal span between checkpoints:
$t_{\text{confirm}}\uparrow\;\Rightarrow\;\text{span}\uparrow,\qquad t_{\text{diagnose}}\downarrow\;\Rightarrow\;\text{span}\uparrow,\qquad t_{\text{redo}}\downarrow\;\Rightarrow\;\text{span}\uparrow.$
C.2. Sensitivity to Per-Step Reliability
We conducted a simple sensitivity analysis by sweeping the per-step correctness rate $p$ from 0.60 to 0.99 and computing the corresponding optimal checkpoint frequency. As shown in Fig. 13, in this 16-step task, the number of confirmation points decreases sharply as reliability increases. When $p\leq 0.70$, the model places checkpoints at nearly every step, as the compounding error risk quickly overwhelms the cost of confirmation. Once accuracy exceeds $p=0.80$, checkpoint frequency begins to drop rapidly, and at near-perfect accuracy ($p=0.99$), the schedule collapses to a single confirmation at the end.
Figure 13. Sensitivity of the optimal checkpoint schedule to uniform per-step correctness in the image-editing domain.
A bar chart where the x-axis is per-step correctness and the y-axis is the number of checkpoints in the optimal schedule for a 16-step image-editing task. Bars are higher for lower \$p\$, showing that the scheduler responds to reduced reliability by increasing confirmation frequency.
Appendix D Simulation on Probability-Based Baselines
Figure 14. Shopping: Estimated completion time under probability-based confirmation strategies (green) versus our model’s scheduler (red).
Bar chart for the shopping domain with expected total time on the y-axis and per-step confirmation probability on the x-axis. Several light-green bars show different probability levels; the red “Our model” bar is close to the lowest green bar, illustrating that the model’s schedule is as good as or slightly better than the best probability-based strategy.
Figure 15. Image-editing: Estimated completion time under probability-based confirmation strategies (green) versus our model’s scheduler (red).
Bar chart for the image-editing domain with expected total time on the y-axis and per-step confirmation probability on the x-axis. Light-green bars show the expected time for different confirmation probabilities, forming a shallow U-shape. A red bar labeled “Our model” appears near the bottom of the U, indicating that the model’s scheduler performs as well as or better than any of the probability-based baselines.
Figure 16. Overcooked: Estimated completion time under probability-based confirmation strategies (green) versus our model’s scheduler (red).
Bar chart for the Overcooked domain with expected total time on the y-axis and per-step confirmation probability on the x-axis. Light-green bars show the expected time for different confirmation probabilities, forming a shallow U-shape. A red bar labeled “Our model” appears near the bottom of the U, indicating that the model’s scheduler performs as well as or better than any of the probability-based baselines.
As a complement to our user study, we ran a stochastic simulation that instantiates a family of probability-based confirmation policies and compares them against our model’s checkpoint schedule. In all three domains (shopping, image editing, and Overcooked), we reused exactly the same parameters as in the user study, including task length, per-step action accuracy, and the timing parameters for confirmation, diagnosis, and redo.
Under the probability-based policy, each step has an independent probability $p_{\text{check}}$ of becoming a checkpoint. For an $N$-step task, this yields an expected number of checkpoints of approximately $p_{\text{check}}*N$. To cover a wide range of behaviors, we swept $p_{\text{check}}$ over values whose expected number of checkpoints ranges from roughly 1 to $N$, thus spanning almost all reasonable confirmation frequencies.
For each schedule, we simulated 5,000 random runs. In each run, every step independently succeeded or failed according to the per-step accuracy, and we applied the CDCR pattern to compute the total completion time based on when checkpoints occurred and where the first error appeared. We then averaged the completion times over these 5,000 runs to obtain an empirical estimate of the expected time for that schedule.
Figures 14 16 plot the estimated completion time for all probability-based baselines, with an additional bar for our model. Across all three domains, the probability-based policies exhibit the expected U-shaped pattern: very small $p_{\text{check}}$ values lead to few checkpoints and large rollback costs, while very large $p_{\text{check}}$ values approach confirming at every step and incur high confirmation overhead. Our model’s schedule (red bar) consistently lies the bottom of this curve and slightly lower than, the minimum across all probability baselines in each domain. Because our model uses the same number of checkpoints as the best-performing probability baseline, but optimizes their exact placement rather than only fixing the total count.